QUESTION 7: You bought 250 shares of McWeber Inc. stock on margin. Price paid was $75 per share. The initial margin is 60%. The maintenance margin is 25%. A. What is the initial equity per share? (2 points) B. What is the loan amount per share? (2 points) C. How low can the price fall before there is a margin call? (2 points) D. Assume the price falls to $36. What is the amount of the margin call? (3 points) E. Assume the price did not fall. Instead the price increased to $92 and you sold the stock. During the nine month period in which you held the stock you received dividends of $.82 each quarter. Assume also that the annual interest rate charged on the loan amount is 8%. What is the holding period return on your invested capital (i.e. your equity investment) for the 9-month holding period? (5 points)
Answer #1

A. What is the initial equity per share? (2 points)
=Price*initial margin
=75*60%=45

B. What is the loan amount per share? (2 points)
=Price*(1-initial margin)
=75*(1-60%)=30

C. How low can the price fall before there is a margin call? (2 points)
=Loan amount/(1-maintenance margin)
=30/(1-25%)=40

D. Assume the price falls to $36. What is the amount of the margin call? (3 points)
=250*60%*36-(250*75*60%-250*(75-36))
=3900

P.S.: I am not allowed to answer more than 4 questions